OpenStudy (anonymous):

Evaluate the indefinite integral: ln(7x)dx.

OpenStudy (whynot):

(x(log(7x)-1)+C

OpenStudy (anonymous):

it can be reduced to integral(t*e^t/7 dt) and apply integration of products rule to get (t-1)e^t/7+Constant of integration.

OpenStudy (amistre64):

i tend to see it as int by parts

OpenStudy (anonymous):

it is int by parts

OpenStudy (whynot):

it is by parts

OpenStudy (amistre64):

\[\int ln(7x)dx=x\ ln(7x)-\int {\frac{7x^2}{27x}} dx\] \[\int ln(7x)dx=x\ ln(7x)-\int {\frac{x}{2}} dx\] maybe

OpenStudy (amistre64):

i misplaced a few things lol

myininaya (myininaya):

\[\int\limits_{}^{}(\ln(7)+\ln(x)) dx=\ln(7) \cdot x+(x \ln(x)-\int\limits_{}^{}1 dx)+C=\ln(7) \cdot x+x \ln(x)-x+C\]

OpenStudy (amistre64):

\[\int ln(7x)dx=x(v)\ ln(7x)(u)-\int {x(v)\frac{1}{x}(du)} dx\] \[\int ln(7x)dx=x\ ln(7x)-\int {x\frac{1}{x}} dx\] \[\int ln(7x)dx=x\ ln(7x)-\int dx\]

OpenStudy (amistre64):

\[\int ln(7x)dx=x\ ln(7x)-x+C\] dbl chk with a derivative of it tho

myininaya (myininaya):

looks good amistre but i think it is easier to write ln(7x) as ln(7)+ln(x) then integrate

OpenStudy (amistre64):

your right, but i tend to do my math like a do my women :) in the end, aint nobody happy

myininaya (myininaya):

lol